OpenStudy (anonymous):

Calculate the effective rate (APY) of interest for 1 year. The principal is $15,500 with an interest rate of 12%. The interest is compounded quarterly.

OpenStudy (anonymous):

The interest is compounded quarterly... You know that quarterly means 4 times a year. It's 12% per year, so that means if you slice that into fourths, you'll get 3 for every 4. Finish the problem as there's a "3" for every 1/4 of a year. So that transforms 1 to 1.03. So we have four times a year! 1.03^4 = 1.12550881. Subtract 100% (or 1) from that, and you're left with 0.12550881. What is that in percentage? 12.55%!

OpenStudy (anonymous):

(1+i) = (1 + 0.12/4)^4 1+i = 1.12550881 i = .12550881 i = .12550881x 100 i = 12.5%